OpenStudy (anonymous):

which line contains the point (0,2) and is perpendicular to y=1/4x+2 ? the choices are : a. y=1/4x+1 b. y=-1/4x+2 c. y=-4x+2 d. y=4x+2

OpenStudy (anonymous):

first off, we need to find the slope of a perpendicular line from the line of y=1/4x+2 if m1 = 1/4 find what m2 would be in: m1*m2=-1

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

mm... how a bout like this: m*(1/4) = -1 solve for m

OpenStudy (anonymous):

is it -4 ?

OpenStudy (anonymous):

yep so the slope, m, for our perpendicular line to the equation y=1/4x+2, is -4

OpenStudy (anonymous):

next we take the point given (0, 2) and m=-4 substitue those into y = mx + b then solve for b

OpenStudy (anonymous):

2=-4(0)+b 2=b is that right ?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

so, the next thing then is to substittue in m and b into y=mx+b and compare it with the choices given in the problem

OpenStudy (anonymous):

ok got it thanks :*

OpenStudy (anonymous):

^_^